# ReCon: Reconfiguring Analog Rydberg Atom Quantum Computers for Quantum Generative Adversarial Networks

### v1

@@ 1. Introduction
The Opportunity Gap. While GANs can theoretically benefit from quantum computing, current noisy intermediate-scale quantum (NISQ) computers suffer from a variety of challenges related to hardware errors and constraints (Patel and Tiwari, 2020; Henriet et al., 2020; Patel et al., 2020a). For instance, prior quantum GANs have been implemented on the superconducting-qubit technology, which has benefited from rapid development due to being built on readily-available silicon subtrates (Li et al., 2020; Arute et al., 2019; Smith et al., 2022). However, this technology has challenges related to short qubit decoherence times, lack of operations involving more than two qubits, and hardwired qubit connectivity (Burnett et al., 2019; Li et al., 2020; Patel et al., 2020b). While the recently demonstrated Rydberg atom quantum computers also suffer from hardware errors, it has advantages such as longer qubit decoherence times, multi-qubit operations, and reconfigurable qubit connectivity (Henriet et al., 2020; Saffman, 2019; Tan et al., 2022).
@@ 2. Relevant Background
This representation generalizes to systems having higher qubit counts. We combine the individual Hilbert spaces of qubits via the tensor product operation. For two qubits, this results in a combined 4-dimensional space that is spanned by basis vectors $|00\rangle$, $|01\rangle$, $|10\rangle$, and $|11\rangle$. We can extend this idea to more than two qubits: for example, a 3-qubit system will have an 8-dimensional Hilbert space spanned by $|000\rangle$, $|001\rangle$, $|010\rangle$, $|011\rangle$, $|100\rangle$, $|101\rangle$, $|110\rangle$, and $|111\rangle$. In general, $n$ qubits will have a combined state |$\ket{\Psi}=\sum_{k=0}^{k=2^{n}-1}\alpha_{k}\ket{k}$, such that $\sum_{k=0}^{k=2^{n}-1}\norm{\alpha_{k}}^{2}=1$. The probability of observing the $k^{th}$ state is $\norm{\alpha_{k}}^{2}$. The Hilbert space size for an $n-$qubit system is $2^{n}$.
@@
Rydberg Atom Quantum Computers. In a Rydberg atom computer, qubits are stored as the electronic energy levels of neutral atoms (Henriet et al., 2020; Wurtz et al., 2023). Generally, these computers use a low energy ground state, |$\ket{g}$, and a highly excited state |$\ket{r}$, referred to as a Rydberg state. One possible scheme uses these two states as our qubit states upon defining |$\ket{0}=\ket{g}$ and |$\ket{1}=\ket{r}$. We will use these notations interchangeably. We also use the terms ”qubits“ and ”atoms“ interchangeably, as they represent the same entity in this context.
@@
Here, $|\Psi(t)$ represents the state of the quantum system at time $t$, and $\hbar$ is the reduced Planck’s constant. Unlike digital quantum computers, which define program execution in terms of discrete quantum gates, analog systems execute quantum programs by manipulating the system Hamiltonian as the state evolves in time, as shown in Fig. 1b(a). In general, Hamiltonians will have different controllable parameters, depending explicitly on the type of quantum computer used. In Rydberg atom computers, these parameters are related to the atoms’ interaction with the laser and the inter-atomic spacing that controls the Rydberg interaction strength.
Advantages of Rydberg Atom Quantum Computers. Highly tunable Rydberg atom quantum computers provide the flexibility to finely tune the interactions between qubits to achieve optimized quantum operations. The Rydberg interaction depends inversely on the separation between atoms, so as atoms are brought closer together, they interact more strongly with one another. As demonstrated in the first configuration topology in Fig. 1b(b), Q0 and Q1 have the strongest dipole-dipole interaction, while Q2 and Q3 have the weakest. This allows the computer to strongly couple Q0 and Q1 so that their states may be entangled. The exact positioning will have important ramifications for how correlations form between qubits in different parts of the computer.
@@
Generative Adversarial Networks(GANs). GANs consist of two networks: a generator and a discriminator (Fig. 3). In a conventional image generation setting, the generator is intended to learn such that it produces synthetic images that are indistinguishable from real images. Ideally, the generator will produce a variety of high-quality images. To introduce this variability, the generator takes in seed noise, which is typically drawn from some prior distribution. The generator then maps this seed noise $Z$ to an output image $G$ by way of parameterized functions/layers:
(2) $G(Z;\vec{\theta})=(g_{0},g_{1},...,g_{k})$
Here, the $g_{i}^{\prime}s$ are output features (pixels in the case of image generation). The parameters $\vec{\theta}$ are then optimized so that the generator produces high-quality images across a variety of input seed noise. This is done through adversarial training against the discriminator network, which is tasked with identifying which images are real and which images are synthetic. As the discriminator network improves in being able to distinguish between real and synthetic input images, the generator network parameters are updated by minimizing the probability that its output belongs to the synthetic class, as assigned by the discriminator network. The learning process of a GAN can thus be summarized as a back-and-forth competition between the two networks. A successfully developed generator network should be able to produce diverse, realistic images.
@@ 4. Design and Implementation
$g_{i}=p_{i}\ \text{mod}\ \frac{1}{2^{n}}$
Now, each generator output $g_{i}$ can be any value on the interval $(0,\frac{1}{2^{n}}]$, regardless of the other outputs. While training, we scale the training data PCA features so they are also in this allowed range. Following training, we rescale all generated PCA features before applying the PCA inverse transform.
Generator Hamiltonian Parameters. In Rydberg atom quantum computers, laser pulses are used to manipulate the Hamiltonian, with users being able to control the pulses via four parameters: the Rabi frequency ($\Omega$), the detuning $(\Delta)$, the phase ($\phi$), and the atoms’ positions within the atom grid ($\vec{x}$). The Rabi frequency is a measure of the intensity of the laser pulse that drives the transitions between the ground state and the excited Rydberg states of qubits. This parameter controls the rate at which qubits oscillate between these two states. The detuning, $(\Delta)$, refers to the difference between the frequency of the laser and the natural frequency of the atomic transition. Finally, the relative phase $\phi$ controls the direction in which the pulse rotates each qubit’s state (e.g., in the X or Y directions). By controlling these parameters over a length of time, user-defined evolutions in the Hamiltonian can be performed (Henriet et al., 2020; Parra-Rodriguez et al., 2020; Wurtz et al., 2023).
@@
(3) $\begin{split}H(t)=\frac{\Omega(t)}{2}\sum_{i}\left(e^{i\phi(t)}|g\rangle_{i}\langle r|+e^{-i\phi(t)}|r\rangle_{i}\langle g|\right)\\ -\Delta_{global}(t)\sum_{i}\hat{n}_{i}+\sum_{i<j}\frac{C_{6}}{|\vec{s}_{i}-\vec{s}_{j}|^{6}}\hat{n}_{i}\hat{n}_{j}\end{split}$
In the above, $t$ represents time, and $\Omega(t)$, $\Delta_{global}(t)$, and $\phi(t)$, represent the globally addressed Rabi frequency, detuning, and relative phase at time $t$. Here, global addressing refers to the fact that these Hamiltonian parameters are the same across all atoms. $|g\rangle_{i}$ and $|r\rangle_{i}$ are the ground and excited Rydberg states of atom $i$ respectively ( $\bra{r}=\ket{r}^{\dagger}$, where $\dagger$ represents the Hermitian conjugate). $\hat{n}_{i}$ counts the number of Rydberg excitations for atom $i$ and is equivalent to $|r\rangle_{i}\langle r|_{i}$. $\frac{C_{6}}{|\vec{s}_{i}-\vec{s}_{j}|^{6}}$ represents the Rydberg interaction potential, with $\vec{x}_{i}$ being the location of the atom $i$ in the grid and the constant $C_{6}=5,420,503$ MHz $\mu m^{6}$.
@@
ReCon encodes both the generator’s input noise and trainable parameters into the laser pulse sequence and positions. For the former, this requires specifying a time series that depends non-trivially on both the input noise and the generator’s trainable parameters.
@@
Pulse Parameterization. We design ReCon to use a selection of simple functions as possible pulse shapes, shown in Fig. 7. Doing so allows ReCon to explore a variety of different Hamiltonian evolutions without having to rely on a diverse parameterization scheme. Each pulse shape depends on a single scalar parameter, which is trainable, and the input noise is used as a seed for image generation. As a concrete example, consider the linear pulse shape of Fig. 7. One possible parameterization is to set the initial value of the pulse based on the seed noise and then the final value based on the pulse parameter, with intermediate values determined by linear interpolation. The goal of training is then to find a suitable pulse parameter that allows for a variety of high-quality images to be made from different initial seeds.
@@
$ValData\leftarrow\text{Validation set of real data}$
$ValFID\leftarrow$ 0
[…]
@@
         if $TrialFID<ValFID\text{{ or }}Ensemble=\{\text{\O}\}$ then
              $Ensemble$.append(L)
@@ 5. Methodology
Simulation with Hardware Errors. To simulate hardware errors, ReCon follows a similar model to the one that is used in (Lu et al., 2024), which is based on errors from the Aquila Rydberg computer (Wurtz et al., 2023). Generally, errors are modeled as a Gaussian perturbation added to the ideal parameters. The detuning and Rabi frequency experience a stochastic shift modeled as $\Delta+N(0,0.1MHz)$ (both local and global detuning follow this error model) and $\Omega+N(0,0.01)$, respectively, while the coordinates of atom $j$ are perturbed by a Gaussian process resulting in $x_{j}+N(0,0.01\mu m)$. Here, $N(\mu,\sigma)$ is a Gaussian distribution with $\mu$ being the mean and the $\sigma$ being the standard deviation.
Quantum Hardware. For real hardware simulation, we use QuEra’s Aquila device through the Amazon Braket interface (Gonzalez, 2021). At the time of writing, Aquila is readily capable of global addressing for the Rabi frequency and detuning and features an experimental implementation of locally addressable detuning. We use 1000 shots to get an empirical estimate of the probability distribution. The modulo operation and inverse PCA transform are then applied during postprocessing to convert the quantum output into images. For hardware compatibility, we must discretize the pulses so that they can be converted into an instruction set for the laser. For some pulses, like ”triangle“ and ”trapezoid“, this is not an issue, as they are already piecewise linear functions. For others, like the Gaussian pulse, this discretization will result in a piecewise linear approximation of the pulse shape, which limits our simulation accuracy.

### v2

@@ 1. Introduction
The Opportunity Gap. While GANs can theoretically benefit from quantum computing, current noisy intermediate-scale quantum (NISQ) computers suffer from a variety of challenges related to hardware errors and constraints (Patel and Tiwari, 2020; Henriet et al., 2020; Patel et al., 2020a). For instance, prior quantum GANs have been implemented on superconducting-qubit technology, which has benefited from rapid development due to being built on readily-available silicon subtrates (Li et al., 2020; Arute et al., 2019; Smith et al., 2022). However, this technology has challenges related to short qubit coherence times, lack of operations involving more than two qubits, and hardwired qubit connectivity (Burnett et al., 2019; Li et al., 2020; Patel et al., 2020b). While the recently demonstrated Rydberg atom quantum computers also suffer from hardware errors, they have advantages such as longer qubit coherence times, multi-qubit operations, and reconfigurable qubit connectivity (Henriet et al., 2020; Saffman, 2019; Tan et al., 2022).
@@ 2. Relevant Background
This representation generalizes to systems having higher qubit counts. We combine the individual Hilbert spaces of qubits via the tensor product operation. For two qubits, this results in a combined 4-dimensional space that is spanned by basis vectors $|00\rangle$, $|01\rangle$, $|10\rangle$, and $|11\rangle$. We can extend this idea to more than two qubits: for example, a 3-qubit system will have an 8-dimensional Hilbert space spanned by $|000\rangle$, $|001\rangle$, $|010\rangle$, $|011\rangle$, $|100\rangle$, $|101\rangle$, $|110\rangle$, and $|111\rangle$. In general, $n$ qubits will have a combined state |$\ket{\Psi}=\sum_{k=0}^{k=2^{n}-1}\alpha_{k}\ket{k}$, such that $\sum_{k=0}^{k=2^{n}-1}\absolutevalue{\alpha_{k}}^{2}=1$. The probability of observing the $k^{th}$ state is |$\absolutevalue{\alpha_{k}}^{2}$. The Hilbert space size for an $n-$qubit system is $2^{n}$.
@@
Rydberg Atom Quantum Computers. In a Rydberg atom computer, qubits are stored as the electronic energy levels of neutral atoms (Henriet et al., 2020; Wurtz et al., 2023). Generally, these computers use a low energy ground state, |$\ket{g}$, and a highly excited state |$\ket{r}$, referred to as a Rydberg state. One possible scheme uses these two states as our qubit states upon defining |$\ket{0}=\ket{g}$ and |$\ket{1}=\ket{r}$. We will use these notations interchangeably. We also use the terms “qubits” and “atoms” interchangeably, as they represent the same entity in this context.
@@
Here, $|\Psi(t)\rangle$ represents the state of the quantum system at time $t$, and $\hbar$ is the reduced Planck’s constant. Unlike digital quantum computers, which define program execution in terms of discrete quantum gates, analog systems execute quantum programs by manipulating the system Hamiltonian as the state evolves in time, as shown in Fig. 1b(a). In general, Hamiltonians will have different controllable parameters, depending explicitly on the type of quantum computer used. In Rydberg atom computers, these parameters are related to the atoms’ interaction with the laser and the inter-atomic spacing that controls the Rydberg interaction strength.
Advantages of Rydberg Atom Quantum Computers. Highly tunable Rydberg atom quantum computers provide the flexibility to finely tune the interactions between qubits to achieve optimized quantum operations. The Rydberg interaction depends inversely on the separation between atoms, so as atoms are brought closer together, they interact more strongly with one another. As demonstrated in the first configuration topology in Fig. 1b(b), Q0 and Q1 have the strongest dipole-dipole interaction, while Q1 and Q2 have the weakest. This allows the computer to strongly couple Q0 and Q1 so that their states may be entangled. The exact positioning will have important ramifications for how correlations form between qubits in different parts of the computer.
@@
Generative Adversarial Networks(GANs). GANs consist of two networks: a generator and a discriminator (Fig. 3). In a conventional image generation setting, the generator is intended to learn such that it produces synthetic images that are indistinguishable from real images. Ideally, the generator will produce a variety of high-quality images. To introduce this variability, the generator takes in seed noise, which is typically drawn from some prior distribution. The generator then maps this seed noise $Z$ to an output image $F$ by way of parameterized functions/layers:
(2) $F(Z;\vec{\theta})=(f_{0},f_{1},...,f_{k})$
Here, the $f_{i}^{\prime}s$ are output features (pixels in the case of image generation). The parameters $\vec{\theta}$ are then optimized so that the generator produces high-quality images across a variety of input seed noise. This is done through adversarial training against the discriminator network, which is tasked with identifying which images are real and which images are synthetic. As the discriminator network improves in being able to distinguish between real and synthetic input images, the generator network parameters are updated by minimizing the probability that its output belongs to the synthetic class, as assigned by the discriminator network. The learning process of a GAN can thus be summarized as a back-and-forth competition between the two networks. A successfully developed generator network should be able to produce diverse, realistic images.
@@ 4. Design and Implementation
$f_{i}=p_{i}\ \text{mod}\ \frac{1}{2^{n}}$
Now, each generator output $f_{i}$ can be any value on the interval $(0,\frac{1}{2^{n}}]$, regardless of the other outputs. While training, we scale the training data PCA features so they are also in this allowed range. Following training, we rescale all generated PCA features before applying the PCA inverse transform.
Generator Hamiltonian Parameters. In Rydberg atom quantum computers, laser pulses are used to manipulate the Hamiltonian, with users being able to control the pulses via three parameters: the Rabi frequency ($\Omega$), the detuning $(\Delta)$, and the phase ($\phi$). The Rabi frequency is a measure of the intensity of the laser pulse that drives the transitions between the ground state and the excited Rydberg states of qubits. This parameter controls the rate at which qubits oscillate between these two states. The detuning, $(\Delta)$, refers to the difference between the frequency of the laser and the natural frequency of the atomic transition. Finally, the relative phase $\phi$ controls the direction in which the pulse rotates each qubit’s state (e.g., in the X or Y directions). By controlling these parameters over a length of time, user-defined evolutions in the Hamiltonian can be performed (Henriet et al., 2020; Parra-Rodriguez et al., 2020; Wurtz et al., 2023).
@@
(3) $\begin{split}H(t)=\frac{\Omega(t)}{2}\sum_{i}\left(e^{i\phi(t)}|g\rangle_{i}\langle r|_{i}+e^{-i\phi(t)}|r\rangle_{i}\langle g|_{i}\right)\\ -\Delta_{global}(t)\sum_{i}\hat{n}_{i}+\sum_{i<j}\frac{C_{6}}{|\vec{s}_{i}-\vec{s}_{j}|^{6}}\hat{n}_{i}\hat{n}_{j}\end{split}$
In the above, $t$ represents time, and $\Omega(t)$, $\Delta_{global}(t)$, and $\phi(t)$, represent the globally addressed Rabi frequency, detuning, and relative phase at time $t$. Here, global addressing refers to the fact that these Hamiltonian parameters are the same across all atoms. $|g\rangle_{i}$ and $|r\rangle_{i}$ are the ground and excited Rydberg states of atom $i$ respectively ( $\bra{r}=\ket{r}^{\dagger}$, where $\dagger$ represents the Hermitian conjugate). $\hat{n}_{i}$ counts the number of Rydberg excitations for atom $i$ and is equivalent to $|r\rangle_{i}\langle r|_{i}$. $\frac{C_{6}}{|\vec{s}_{i}-\vec{s}_{j}|^{6}}$ represents the Rydberg interaction potential, with $\vec{s}_{i}$ being the location of the atom $i$ in the grid and the constant $C_{6}=5,420,503$ MHz $\mu m^{6}$.
@@
The parameters of ReCon’s generator are thus the laser pulse sequence, atom positions, and site-dependent local detuning coupling. The atomic positions and detuning coupling are both fixed at the beginning of the analog computation. The pulses are each specified by a time series. To incorporate the generator’s input seed noise, we design pulses that depend non-trivially on some scalar input noise and a set of trainable parameters.
@@
Pulse Parameterization. We design ReCon to use a selection of simple functions as possible pulse shapes, shown in Fig. 7. Doing so allows ReCon to explore a variety of different Hamiltonian evolutions without having to rely on a complex parameterization scheme. Each pulse shape depends on a single scalar parameter, which is trainable, and the input noise is used as a seed for image generation. As a concrete example, consider the linear pulse shape of Fig. 8. One possible parameterization is to set the initial value of the pulse based on the seed noise and then the final value based on a single scalar pulse parameter, with intermediate values determined by linear interpolation. The goal of training is then to find a suitable pulse parameter that allows for a variety of high-quality images to be made from different initial seeds.
@@
Inputs: $ValData$ (validation set of real data), $bestLearner$ (best learner measured against ValData)
$Ensemble\leftarrow\{bestLearner\}$
$ValFID\leftarrow$ FID of $bestLearner$
$Searching\leftarrow True$
@@
         if $TrialFID<ValFID$ then
              $nextLearner\leftarrow L$
@@
     if $nextLearner=nothing$ then
         $Searching\leftarrow False$
     else
         $Ensemble\leftarrow Ensemble\cup\{nextLearner\}$
     end if
@@ 5. Methodology
Simulation with Hardware Errors. To simulate hardware errors, ReCon follows a similar model to the one that is used in (Lu et al., 2024), which is based on errors from the Aquila Rydberg computer (Wurtz et al., 2023). Generally, errors are modeled as a Gaussian perturbation added to the ideal parameters. The detuning and Rabi frequency experience a stochastic shift modeled as $\Delta+N(0,0.1\text{ MHz})$ (both local and global detuning follow this error model) and $\Omega\cdot N(1\,,0.01\text{ MHz})$, respectively, while the coordinates of atom $j$ are perturbed by a Gaussian process resulting in $x_{j}+N(0\,,0.1\text{ $\mu$m})$ (the same for $y_{i}$ as well). Here, $N(\mu,\sigma)$ is a Gaussian distribution with $\mu$ being the mean and the $\sigma$ being the standard deviation.
Quantum Hardware. For real hardware simulation, we use QuEra’s Aquila device through the Amazon Braket interface (Gonzalez, 2021). At the time of writing, Aquila is readily capable of global addressing for the Rabi frequency and detuning and features an experimental implementation of locally addressable detuning. We use 1000 shots to get an empirical estimate of the probability distribution. The modulo operation and inverse PCA transform are then applied during postprocessing to convert the quantum output into images. For hardware compatibility, we must discretize the pulses so that they can be converted into an instruction set for the laser. For some pulses, like “triangle” and “trapezoid”, this is not an issue, as they are already piecewise linear functions. For others, like the Gaussian pulse, this discretization will result in a piecewise linear approximation of the pulse shape, which limits our simulation accuracy.